# DeepResearch Arena: The First Exam of LLMs’ Research Abilities via Seminar-Grounded Tasks

### v1

@@ Introduction
Recent developments in large language models (LLMs) have led to the rise of the deep research agent  (Huang et al. 2025; Xu and Peng 2025; Wu et al. 2025), a LLM-powered agentic system designed for research task automation by integrating literature search (Baek et al. 2024), experiment design (Schmidgall et al. 2025), and ideation (Li et al. 2024). Prevailing examples, such as GPT DeepResearch (OpenAI 2025), indicate that deep research agents have great potential to significantly promote research creativity and productivity.
While deep research agents have gained increasing attention (Du et al. 2025), faithfully evaluating their research ability remains a huge challenge. As Einstein once stated, The formulation of the problem is often more essential than its solution, which may be merely a matter of mathematical or experimental skill (Einstein and Infeld 1938). This perspective highlights a crucial challenge in formulating high-quality and frontier research tasks to faithfully assess the ability of deep research agents.
Existing benchmarks for deep research agents mainly resort to two approaches to acquire research questions. The first leverages static data corpora such as academic literature and web content, as seen in AcademicBrowse (Zhou et al. 2025a), BrowseComp (Wei et al. 2025), and Researchbench (Liu et al. 2025). The second approach involves manually curated research tasks by domain experts, exemplified by Humanity’s Last Exam (Phan et al. 2025), DeepResearchBench (Du et al. 2025), and ExpertLongBench (Ruan et al. 2025). However, both approaches are hindered by critical limitations. Benchmarks derived from static corpora risk data leakage, as the underlying content may already be included in the model pertaining. Meanwhile, datasets curated by experts face scalability bottlenecks and often lack the diversity and spontaneity found in authentic research settings. More fundamentally, both sources tend to abstract away from the situated, evolving nature of real-world research inquiry, where questions emerge dynamically through discourse, ambiguity, and interdisciplinary exploration. A detailed comparison of these benchmarks across key dimensions, including scalability, automation, data leakage risk, and research realism, is provided in Table 1.
@@
To capture the nature of such authentic inquiry, Deep Research Arena formulates tasks as open-ended, under-defined problems, drawn by the theory of Ill-Structured Problem Solving (Jonassen 1997), which describes real-world problems as “poorly defined, with no single correct formulation and no objective evaluation criteria”. To construct the Deep Research Arena, we develop a Multi-Agent Hierarchical Task Generation (MAHTG) system that automatically extracts research-worthy inspirations and systematically transforms them into high-quality, traceable research tasks through a multi-stage filtering and structuring pipeline. This design ensures both the authenticity and reproducibility of task construction, while reducing noise and preserving the intellectual context of original expert discourse.
@@
The emergence of DR agents builds upon recent advances in LLMs equipped with tool-use capabilities (Li et al. 2025; Qu et al. 2025; Tang et al. 2023), which allow models to interface with search engines, code interpreters, and external APIs to extend their reasoning horizon. On this foundation, systems such as GPT Deep Research (OpenAI 2025), Gemini Deep Research (Google 2025), and Grok DeepSearch (xAI 2025) have been developed to support multi-stage research workflows. GPT’s system focuses on outline-driven long-form synthesis with citation grounding, Gemini emphasizes multimodal retrieval and synthesis, while Grok prioritizes real-time web summarization for dynamic topics. These agents reflect a shift from retrieval-based assistants to goal-directed, tool-augmented agents capable of supporting exploratory, open-ended inquiry (Yu, He, and Ying 2023).
@@
Existing benchmarks for deep research agents mainly resort to two approaches to acquire research questions: automatically deriving tasks from static corpora or manually curating them through expert design. The first leverages static data corpora such as papers, and web documents to construct benchmarks represented by multi-hop reasoning or simplified scientific queries. Examples include MuSiQue (Trivedi et al. 2022), which automatically generates multi-hop questions by linking single-hop QA pairs from existing datasets, and HotpotQA (Yang et al. 2018), where annotators write questions guided by system-selected Wikipedia article pairs, making the process closer to extraction than genuine question generation. Other benchmarks in this category include StrategyQA (Geva et al. 2021), ThoughtSource (Ott et al. 2023), AcademicBrowse (Zhou et al. 2025a), and BrowseComp (Wei et al. 2025). Despite their emphasis on multi-step reasoning, these benchmarks rely on manually constructed logic chains with predefined paths. They primarily test factual retrieval and compositional reasoning capabilities, yet fail to capture how research questions naturally emerge, evolve, and iterate in real-world research contexts. ScienceQA (Lu et al. 2022) is a large-scale multimodal multiple-choice science QA benchmark ( 21K questions across STEM and social/language science) that includes lecture and explanation‑level CoT annotations to support interpretable multi-step reasoning.
@@
The second category consists of expert-authored benchmarks, where researchers collaborate with domain specialists to construct high-quality, PhD-level evaluation tasks. Compared to benchmarks built from static corpora, these datasets typically feature more original, conceptually challenging, and discipline-specific questions that better reflect expert-level reasoning. Representative examples include LAB-Bench (Laurent et al. 2024), ARC (Clark et al. 2018), GPQA (Rein et al. 2024), FrontierMath (Glazer et al. 2024), and Humanity’s Last Exam (Phan et al. 2025). GPQA provides graduate-level multiple-choice questions in biology, physics, and chemistry, curated and verified by domain PhDs to ensure they cannot be solved via surface-level heuristics or web search. Humanity’s Last Exam comprises a collection of open-ended, expert-written research questions across disciplines such as history, philosophy, and theoretical science, designed to probe creative, integrative thinking under minimal structural constraints. DeepResearch Bench (Du et al. 2025) moves toward more realistic simulation by requiring long-form research reports across disciplines. However, this entire class of expert-authored benchmarks faces several limitations: their prompts are manually constructed, which restricts scalability and diversity, and the datasets remain relatively small in size. More fundamentally, they also fail to capture how research questions emerge dynamically through discourse, ambiguity, and interdisciplinary exploration—core characteristics of authentic research practice.
@@ III. Task Phase Labels
Building on the structured inspirations extracted from seminar transcripts, we deploy TaskWeaver Agent that aggregates and reorganizes content across multiple inspirations to synthesize a focused set of concrete research tasks distributed across three key phases—Synthesize, Design, and Evaluate, as illustrated in Figure 2. These tasks are constructed by identifying the core problem focus or methodological cues within the inspirations and are paired with clearly defined, executable goals. This structured synthesis process enables the scalable construction of diverse, high-quality DeepResearch tasks aligned with the demands of real-world scientific inquiry (Yu et al. 2022).
To rank the quality of research tasks, we adopt RankEval Agent based on the Elo rating system (Glickman 1995). Each task is initialized with a rating score of 1200. In each round, we randomly sample disjoint pairs of tasks and compare them based on evaluation criteria such as originality, clarity, and scientific relevance. Given a pair of tasks $t_{a}$ and $t_{b}$ with current Elo scores $r_{a}$ and $r_{b}$, we first compute the expected winning probabilities using:
@@ Appendix A Data Leakage Detection
To verify that our benchmark minimize the risk of data leakage from the pretraining corpora of LLMs, we conduct a comprehensive leakage simulation experiment (Xu et al. 2024; Zhou et al. 2025b) across all 8 evaluated models. This procedure estimates whether any model can reproduce the withheld portion of a task when prompted with only the first half of the task description.
@@ A.2 Similarity Metrics
We compute string-level similarity between the model-generated suffix and the ground-truth suffix using the normalized Longest Common Subsequence (LCS) metric (Öztürk, Ozkirimli, and Özgür 2016; Wang and Dong 2020). The similarity score for model $\mathcal{M}_{k}$ is defined as:
@@
We compute lexical similarity between the generated suffix and the reference suffix using cosine similarity over their TF-IDF representations (Castaño et al. 2016; Wang and Dong 2020). The score for model $\mathcal{M}_{k}$ is given by:
@@
We compute word-level lexical overlap between the generated and reference suffixes using the normalized word set intersection (Jindal and Leema 2024). The score for model $\mathcal{M}_{k}$ is defined as:
@@
To obtain a unified similarity score that balances multiple aspects of textual similarity, we compute a weighted combination of the three individual metrics (De Boom et al. 2015):
@@ A.3 Leakage Criterion
We set the threshold $\tau=0.7$ as a conservative criterion to identify potential cases of memorization or contamination. This choice is supported by prior work on training data extraction (Carlini et al. 2021) and sensitive domain leakage (Lehman et al. 2021), which report that similarity scores at or above 0.7 often correspond to memorized or verbatim training content. Such a threshold ensures high precision in detecting potential data leakage while minimizing false positives.
@@ B.2 Experimental Setup
  • Spearman’s Rank Correlation ($\rho$): Measures the monotonic relationship between two ranked variables  (Spearman 1904). It is computed as:
@@
  • Pearson Correlation ($r$): Measures the linear correlation between two variables $X$ and $Y$ (Pearson 1895):
@@
  • Kendall’s Tau ($\tau$): Measures ordinal association by comparing concordant and discordant pairs (Kendall 1938):

### v2

@@ Introduction
Recent developments in large language models (LLMs) have led to the rise of the deep research agent  (huang2025deepresearchagentssystematic; xu2025comprehensivesurveydeepresearch; wu2025agenticreasoningstreamlinedframework), a LLM-powered agentic system designed for research task automation by integrating literature search (baek2024researchagent), experiment design (schmidgall2025agentlaboratoryusingllm), and ideation (li2024chain). Prevailing examples, such as GPT DeepResearch (openai2025deepresearch), indicate that deep research agents have great potential to significantly promote research creativity and productivity.
While deep research agents have gained increasing attention (du2025deepresearch), faithfully evaluating their research ability remains a huge challenge. As Einstein once stated, The formulation of the problem is often more essential than its solution, which may be merely a matter of mathematical or experimental skill (einstein1938evolution). This perspective highlights a crucial challenge in formulating high-quality and frontier research tasks to faithfully assess the ability of deep research agents.
Existing benchmarks for deep research agents mainly resort to two approaches to acquire research questions. The first leverages static data corpora such as academic literature and web content, as seen in AcademicBrowse (zhou2025academicbrowse), BrowseComp (wei2025browsecomp), and Researchbench (liu2025researchbench). The second approach involves manually curated research tasks by domain experts, exemplified by Humanity’s Last Exam (phan2025humanitysexam), DeepResearchBench (du2025deepresearch), and ExpertLongBench (ruan2025expertlongbench). However, both approaches are hindered by critical limitations. Benchmarks derived from static corpora risk data leakage, as the underlying content may already be included in the model pertaining. Meanwhile, datasets curated by experts face scalability bottlenecks and often lack the diversity and spontaneity found in authentic research settings. More fundamentally, both sources tend to abstract away from the situated, evolving nature of real-world research inquiry, where questions emerge dynamically through discourse, ambiguity, and interdisciplinary exploration. A detailed comparison of these benchmarks across key dimensions, including scalability, automation, data leakage risk, and research realism, is provided in Table 1.
@@
To capture the nature of such authentic inquiry, Deep Research Arena formulates tasks as open-ended, under-defined problems, drawn by the theory of Ill-Structured Problem Solving (jonassen1997isp), which describes real-world problems as “poorly defined, with no single correct formulation and no objective evaluation criteria”. To construct the Deep Research Arena, we develop a Multi-Agent Hierarchical Task Generation (MAHTG) system that automatically extracts research-worthy inspirations and systematically transforms them into high-quality, traceable research tasks through a multi-stage filtering and structuring pipeline. This design ensures both the authenticity and reproducibility of task construction, while reducing noise and preserving the intellectual context of original expert discourse.
@@
The emergence of DR agents builds upon recent advances in LLMs equipped with tool-use capabilities (li2025adaptivetooluselarge; Qu_2025; tang2023toolalpacageneralizedtoollearning), which allow models to interface with search engines, code interpreters, and external APIs to extend their reasoning horizon. On this foundation, systems such as GPT Deep Research (openai2025deepresearch), Gemini Deep Research (google2025deepresearch), and Grok DeepSearch (xai2025grok3) have been developed to support multi-stage research workflows. GPT’s system focuses on outline-driven long-form synthesis with citation grounding, Gemini emphasizes multimodal retrieval and synthesis, while Grok prioritizes real-time web summarization for dynamic topics. These agents reflect a shift from retrieval-based assistants to goal-directed, tool-augmented agents capable of supporting exploratory, open-ended inquiry (yu2023thought).
@@
Existing benchmarks for deep research agents mainly resort to two approaches to acquire research questions: automatically deriving tasks from static corpora or manually curating them through expert design. The first leverages static data corpora such as papers, and web documents to construct benchmarks represented by multi-hop reasoning or simplified scientific queries. Examples include MuSiQue (trivedi2022musiquemultihopquestionssinglehop), which automatically generates multi-hop questions by linking single-hop QA pairs from existing datasets, and HotpotQA (yang2018hotpotqa), where annotators write questions guided by system-selected Wikipedia article pairs, making the process closer to extraction than genuine question generation. Other benchmarks in this category include StrategyQA (geva-etal-2021-aristotle), ThoughtSource (Ott_2023), AcademicBrowse (zhou2025academicbrowse), and BrowseComp (wei2025browsecomp). Despite their emphasis on multi-step reasoning, these benchmarks rely on manually constructed logic chains with predefined paths. They primarily test factual retrieval and compositional reasoning capabilities, yet fail to capture how research questions naturally emerge, evolve, and iterate in real-world research contexts. ScienceQA (lu2022learnexplainmultimodalreasoning) is a large-scale multimodal multiple-choice science QA benchmark ( 21K questions across STEM and social/language science) that includes lecture and explanation‑level CoT annotations to support interpretable multi-step reasoning.
@@
The second category consists of expert-authored benchmarks, where researchers collaborate with domain specialists to construct high-quality, PhD-level evaluation tasks. Compared to benchmarks built from static corpora, these datasets typically feature more original, conceptually challenging, and discipline-specific questions that better reflect expert-level reasoning. Representative examples include LAB-Bench (laurent2024labbenchmeasuringcapabilitieslanguage), ARC (clark2018think), GPQA (rein2024gpqa), FrontierMath (glazer2024frontiermathbenchmarkevaluatingadvanced), and Humanity’s Last Exam (phan2025humanitysexam). GPQA provides graduate-level multiple-choice questions in biology, physics, and chemistry, curated and verified by domain PhDs to ensure they cannot be solved via surface-level heuristics or web search. Humanity’s Last Exam comprises a collection of open-ended, expert-written research questions across disciplines such as history, philosophy, and theoretical science, designed to probe creative, integrative thinking under minimal structural constraints. DeepResearch Bench (du2025deepresearch) moves toward more realistic simulation by requiring long-form research reports across disciplines. However, this entire class of expert-authored benchmarks faces several limitations: their prompts are manually constructed, which restricts scalability and diversity, and the datasets remain relatively small in size. More fundamentally, they also fail to capture how research questions emerge dynamically through discourse, ambiguity, and interdisciplinary exploration—core characteristics of authentic research practice.
@@ III. Task Phase Labels
Building on the structured inspirations extracted from seminar transcripts, we deploy TaskWeaver Agent that aggregates and reorganizes content across multiple inspirations to synthesize a focused set of concrete research tasks distributed across three key phases—Synthesize, Design, and Evaluate, as illustrated in Figure 2. These tasks are constructed by identifying the core problem focus or methodological cues within the inspirations and are paired with clearly defined, executable goals. This structured synthesis process enables the scalable construction of diverse, high-quality DeepResearch tasks aligned with the demands of real-world scientific inquiry (yu2022structure).
To rank the quality of research tasks, we adopt RankEval Agent based on the Elo rating system (glickman1995comprehensive). Each task is initialized with a rating score of 1200. In each round, we randomly sample disjoint pairs of tasks and compare them based on evaluation criteria such as originality, clarity, and scientific relevance. Given a pair of tasks $t_{a}$ and $t_{b}$ with current Elo scores $r_{a}$ and $r_{b}$, we first compute the expected winning probabilities using:
@@ Appendix A Data Leakage Detection
To verify that our benchmark minimize the risk of data leakage from the pretraining corpora of LLMs, we conduct a comprehensive leakage simulation experiment (xu2024benchmarkingbenchmarkleakagelarge; zhou2025lessleakbenchinvestigationdataleakage) across all 8 evaluated models. This procedure estimates whether any model can reproduce the withheld portion of a task when prompted with only the first half of the task description.
@@ A.2 Similarity Metrics
We compute string-level similarity between the model-generated suffix and the ground-truth suffix using the normalized Longest Common Subsequence (LCS) metric (ozturk2016comparative; info11090421). The similarity score for model $\mathcal{M}_{k}$ is defined as:
@@
We compute lexical similarity between the generated suffix and the reference suffix using cosine similarity over their TF-IDF representations (castano-etal-2016-machine; info11090421). The score for model $\mathcal{M}_{k}$ is given by:
@@
We compute word-level lexical overlap between the generated and reference suffixes using the normalized word set intersection (jindal2024survey). The score for model $\mathcal{M}_{k}$ is defined as:
@@
To obtain a unified similarity score that balances multiple aspects of textual similarity, we compute a weighted combination of the three individual metrics (De_Boom_2015):
@@ A.3 Leakage Criterion
We set the threshold $\tau=0.7$ as a conservative criterion to identify potential cases of memorization or contamination. This choice is supported by prior work on training data extraction (carlini2021extracting) and sensitive domain leakage (lehman2021does), which report that similarity scores at or above 0.7 often correspond to memorized or verbatim training content. Such a threshold ensures high precision in detecting potential data leakage while minimizing false positives.
@@ B.2 Experimental Setup
  • Spearman’s Rank Correlation ($\rho$): Measures the monotonic relationship between two ranked variables  (spearman1904association). It is computed as:
@@
  • Pearson Correlation ($r$): Measures the linear correlation between two variables $X$ and $Y$ (pearson1895regression):
@@
  • Kendall’s Tau ($\tau$): Measures ordinal association by comparing concordant and discordant pairs (kendall1938rank):